OpenStudy (anonymous):

Evaluate: [(51 + 3) - 3^2] ÷ 9 · 2

OpenStudy (anonymous):

bodmas division then add or subtract

OpenStudy (anonymous):

Using PEMDAS pls

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

The 2 in 32 is an exponent

OpenStudy (anonymous):

first 14/7=2 32+2=34 34-4 30

OpenStudy (texaschic101):

4 - 14/7 + 3^2.....is this the correct problem ?

OpenStudy (anonymous):

yes it is the right problem but i need the answer

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

the answer choices are 11 8 -7 -61

OpenStudy (texaschic101):

first do the exponent...3^2 = 9 then the division....14/7 = 2 now we have 4 - 2 + 9 = 2 + 9 = 11

OpenStudy (anonymous):

thnxs. alot How About this one Evaluate: [(51 + 3) - 3^2] ÷ 9 · 2

OpenStudy (anonymous):

I Really Need Help On This Please.

OpenStudy (texaschic101):

[(51 + 3) - 3^2] /9 *2 Do what is in the brackets first (51 + 3) - 9 /9*2 54 - 9 / 9 *2 45/9 * 2 5 x 2 = 10 You might want to get a second opinion on this.....I am not sure if this is correct

OpenStudy (johnweldon1993):

You are correct @texaschic101 :)

OpenStudy (texaschic101):

I got confused at the end....I didn't know if I should 45/9 x 2.....divide by 9 then multiply by 2 or divide 45 by 18

OpenStudy (johnweldon1993):

When you only have multiplication and division operators left...you just go from left to right *though pemdas says otherwise*

OpenStudy (texaschic101):

thanks for explaining that to me.....thats good to know :)

OpenStudy (johnweldon1993):

of course :)